OpenStudy (anonymous):

For rational function f(x) = (x^2 - 81)/(x^3 - 9x), identify any vertical asymptotes

OpenStudy (anonymous):

To find vertical asymptotes of rational functions, set the denominator equal to zero, and solve.

OpenStudy (anonymous):

so your denominator is \[x^3-9x=0\] take a x out. \[x(x^2-9)=0\] factor \[x(x+3)(x-3)=0\] so what would the vertical asymptotes be?

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

that is one of them, what are the other two?

OpenStudy (anonymous):

x = 3 and x=-3

OpenStudy (anonymous):

thats right